OpenStudy (anonymous):

simplify your answer SR of -20 answer must be typed in form of a+bi

OpenStudy (anonymous):

-20=4*5*-1 2iSR(5)\[2i \sqrt{5}\]

OpenStudy (anonymous):

\[\sqrt{-20}=\sqrt{-1*4*5}=\sqrt{-1}*\sqrt{4}*\sqrt{5}=i*2*\sqrt{5}\]